# Compositional Generative Modeling: A Single Model is Not All You Need

### v1

@@ Abstract
Large monolithic generative models trained on massive amounts of data have become an increasingly dominant approach in AI research. In this paper, we argue that we should instead construct large generative systems by composing smaller generative models together. We show how such a compositional generative approach enables us to learn distributions in a more data-efficient manner, enabling generalization to parts of the data distribution unseen at training time. We further show how this enables us to program and construct new generative models for tasks completely unseen at training. Finally, we show that in many cases, we can discover separate compositional components from data.
@@ 1 Introduction
Despite significant scaling in generative models, existing models remain far from intelligent, exhibiting poor reasoning ability (Tamkin et al., 2021), extensive hallucinations (Zhang et al., 2023), and poor understanding of commonsense relationships in images (Figure 2(Majumdar et al., 2023). Despite this, large models have already been trained on most of the existing data on the Internet and have reached the limits of modern computational hardware, costing hundreds of millions of dollars to train (Figure 1). Inference costs of such gigantic models are also prohibitive, requiring large computational clusters and a cost of several dollars for longer queries and answers (OpenAI, ).
@@
Such compositional generative modeling enables us to effectively represent the natural sparsity and symmetry naturally found in nature. Sparsity of interactions, for instance between an agent and external environment dynamics can be encoded by representing each with separate generative models. Sources of symmetry can be captured using multiple instances of the same independent generative component to represent each occurrence of the symmetry, for instance by tiling patch-level generative model over the patches in an image. Compositional structure is widely used in existing work, to tractably represent high dimensional distributions in Probablistic Graphical Models (PGMs)  (Koller & Friedman, 2009), and even in existing generative models, i.e. autoregressive models which factorize distributions into a set of conditional probability distributions (represented by a single model).
@@ 2 Data Efficient Generative Modeling
The predominant paradigm for training generative models has been to construct increasingly larger monolithic models trained with greater amounts of data and computational power. While language models have demonstrated significant improvements with increased scale (albeit still with difficulty in compositionality (Dziri et al., 2023)), current multimodal models such as DALL-E 3 and GPT-4V remain unable to take advantage of even simple forms of compositionally (Figure 2). Such models may be unable to accurately generate images given combinations of relations rarely seen in training data, or fail to understand simple spatial relations in images, despite being trained on a very significant portion of the existing Internet.
@@
One approach to significantly reduce the amount of data necessary to learn generative models over complex joint distributions is factorization – if we know that a distribution exhibits an independence structure between variables such as
$p(A,B,C,D)\propto p(A)p(B)p(C,D),$
we can substantially reduce the data requirements by only needing to learn these factors, composing them together to form a more complex distribution. This also enables our learned joint distribution to generalize to unseen combinations of variables so long as each local variable combination is in distribution (illustrated in Figure 3). Even in settings where distributions are not accurately modeled as a product of independent factors, such a factorization can still lead to a better generative model given limited data by reducing the hypothesis space (Murphy, 2022). This key idea of factorizing probability distributions has led to substantial work in probabilistic graphical models (PGMs) (Koller & Friedman, 2009).
Below, we illustrate across three settings how representing a target distribution $p(x)$ in a factorized manner can substantially improve generative modeling performance from a limited amount of data:
@@
Compositional Visual Generation.  Finally, we consider modeling a probability distribution $p(x\mid T)$ in text-to-image synthesis, where $x$ is an image and $T$ is a complex text description. While this distribution is usually characterized by a single generative model, we can factor the generation as a product of distributions given sentences $t_{1}$, $t_{2}$, and $t_{3}$ in the description $T$
@@ 3 Generalization to New Distributions
U-Maze Refer to caption
Medium Refer to caption
Large Refer to caption
@@
corresponding to solving the constraint satisfaction problem. Such an approach enables effective generalization to new problems  (Yang et al., 2023b).
@@
Figure 10: Image Tapestries through Composition. By composing a set of probability distributions defined over different spatial regions in an image, we can construct detailed image tapestries. Figure adapted from (Du et al., 2023).
@@
where each distribution is defined over bounding boxes in an image. In Figure 10, we illustrate the efficacy of this approach for constructing complex images.
@@ 4 Generative Modeling with Learned Compositional Structure
where $\text{Enc}(\cdot)$ is a learned neural encoder with low-dimensional latent output to encourage each component to capture distinct regions of an image. By training models to autoencode images with this likelihood expression, each component distribution $p_{\theta}(x\mid\text{Enc}_{i}(x))$ finds interpretable decomposition of images corresponding to individual objects in a scene as well global factors of variation in the scene such as lighting (Du et al., 2021; Su et al., 2024). In Figure 14, we illustrate how these these discovered components, $p_{\theta}(x\mid z_{1})$ and $p_{\theta}(x\mid z_{2})$ from a model trained on cubes and spheres, $p_{\phi}(x\mid z_{3})$ and $p_{\phi}(x\mid z_{4})$ from a separate model trained on trucks and boots can be composed together to form the distribution
@@
Discovering Object Classes From Image Distributions.  Given a distribution of images $p(x)$ representing images drawn from different classes in Imagenet, we can model the likelihood of the distribution as as composition
$p_{\theta}(x)\propto p_{\phi}(w\mid x)\prod_{i}p_{\theta}^{i}(x)^{w_{i}}.$
In Figure 14, we illustrate that the discovered components in this setting represent each of the original Imagenet classes in the input distribution of images. We further illustrate how these discovered components to be composed together to generate images with multiple classes of objects.
@@ 5.1 Challenges With Sampling from Compositional Distributions
where we need to marginalize over all possible future values of $x_{i+1:T}$. Since this marginalization is different dependent on the value of $x_{i}$, $p_{\text{product}}(x_{i}|x_{0:i-1})$ is not equivalent to $p_{1}(x_{i}|x_{0:i-1})p_{2}(x_{i}|x_{0:i-1})$ and therefore autoregressive factorizations are not directly compositional. Similarly, two learned score functions from diffusion models are not directly composable as they do not correspond to the noisy gradient of the product distribution (Du et al., 2023).
@@ 5.2 Effective Compositional Sampling on Continuous Distributions
There has been a substantial body of recent work on improving learning in EBMs (Du & Mordatch, 2019; Nijkamp et al., 2019; Grathwohl et al., 2019; Du et al., 2020b; Grathwohl et al., 2021) but EBMs still lag behind other generative approaches in efficiency and scalability of training. By leveraging the close connection of diffusion models with EBMs in  (Song & Ermon, 2019) we can also directly implement the compositional operations with EBMs with diffusion models  (Du et al., 2023), which we briefly describe below.
Given a diffusion model representing a distribution $p(x)$, we can interpret the $T$ learned denoising functions $\epsilon(x,t)$ of the diffusion model as representing $T$ separate EBM distributions, $e^{-E(x,t)}$, where $\nabla_{x}{E(x,t)}=\epsilon(x,t)$. This sequence of EBM distributions transition from $e^{-E(x,T)}$ representing the Gaussian distribution $\mathcal{N}(0,1)$ to $e^{-E(x,0)}$ representing the target distribution $p_{i}(x)$. We can draw samples from this sequence of EBMs using annealed importance sampling (Du et al., 2023), where we initialize a sample from Gaussian noise and sequentially run several steps of MCMC on each EBM distribution, starting at $e^{-E(x,T)}$ and ending at $e^{-E(x,0)}$.
This EBM interpretation of diffusion models allows them to be composed using operations such as Equation 2 by applying the operation to each intermediate EBM corresponding to the component diffusion distributions, for instance $e^{-(E_{1}(x,k)+E_{2}(x,k))}$. We can then use an annealed importance sampling procedure on this sequence of composite EBMs. Note that this annealed importance procedure is necessary for accurate compositional sampling – using the reverse diffusion process directly on this composed score does not sample from the composed distribution (Du et al., 2023).
@@ 6 Discussion and Future Directions
Such compositional systems have additional benefits in terms of both buildability and interpretability. As individual models are responsible for independent subsets of data, each model can be built separately and modularly by different institutions. Simultaneously, at execution time, it is significantly easier to understand and monitor the execution of each simpler consitituent model than a single large monolithic model.
In addition, such compositional systems can be more environmentally friendly and easier to deploy than large monolithic models. As individual models are substantially smaller, they can run efficiently using small amounts of computation. Simultaneously, it is more straightforward to deploy separate models across separate computational machines.
@@
Finally, while we have provided a few promising results on applications of compositional generative modeling, there remains much to be explored. Future lines of work in area include (1) How do we pass messages and sample from joint distributions? (2) How can we discover the compositional structure inside a generative system? (3) How can we dynamically modify and discover the correct compositional structure of generative models given distribution shift?

### v3

@@ Abstract
Large monolithic generative models trained on massive amounts of data have become an increasingly dominant approach in AI research. We argue that we should instead construct large generative systems by composing smaller generative models together. We show how such a compositional generative approach enables us to learn distributions in a more data-efficient manner, enabling generalization to parts of the data distribution unseen at training time. We further show how this enables us to program and construct new generative models for tasks completely unseen at training. Finally, we show that in many cases, we can discover compositional components from data.
@@ 1 Introduction
Despite significant scaling in generative models, existing models remain far from intelligent, exhibiting poor reasoning ability (Tamkin et al., 2021), extensive hallucinations (Zhang et al., 2023b), and poor understanding of commonsense relationships in images (Figure 2(Majumdar et al., 2023). Despite this, large models have already been trained on most of the existing data on the Internet and have reached the limits of modern computational hardware, costing hundreds of millions of dollars to train (Figure 1). Inference costs of such gigantic models are also prohibitive, requiring large computational clusters and a cost of several dollars for longer queries and answers (OpenAI, ).
@@
Such compositional generative modeling enables us to effectively represent the sparsity and symmetry naturally found in nature. Sparsity of interactions, for instance between an agent and external environment dynamics can be encoded by representing each with separate generative models. Sources of symmetry can be captured using multiple instances of the same independent generative component to represent each occurrence of the symmetry, for instance by tiling patch-level generative model over the patches in an image. Compositional structure is widely used in existing work, to tractably represent high dimensional distributions in Probablistic Graphical Models (PGMs)  (Koller & Friedman, 2009), and even in existing generative models, i.e. autoregressive models which factorize distributions into a set of conditional probability distributions (represented by a single model).
@@ 2 Data Efficient Generative Modeling
The predominant paradigm for training generative models has been to construct increasingly larger monolithic models trained with greater amounts of data and computational power. While language models have demonstrated significant improvements with increased scale (albeit still with difficulty in compositionality (Dziri et al., 2023)), current multimodal models such as DALL-E 3 and GPT-4V remain unable to take advantage of even simple forms of compositionality (Figure 2). Such models may be unable to accurately generate images given combinations of relations rarely seen in training data, or fail to understand simple spatial relations in images, despite being trained on a very significant portion of the existing Internet.
@@
One approach to significantly reduce the data necessary to learn generative models over complex joint distributions is factorization – if we know that a distribution exhibits an independence structure between variables such as
$p(A,B,C,D)\propto p(A)p(B)p(C,D),$
we can substantially reduce the data requirements by only needing to learn these factors, composing them together to form a more complex distribution. This also enables our learned joint distribution to generalize to unseen combinations of variables so long as each local variable combination is in distribution (illustrated in Figure 3). Even in settings where distributions are not accurately modeled as a product of independent factors, such a factorization can still lead to a better models given limited data by reducing the hypothesis space (Murphy, 2022). This idea of factorizing probability distributions has led to substantial work in probabilistic graphical models (PGMs) (Koller & Friedman, 2009).
Below, we illustrate across four settings how representing a target distribution $p(x)$ in a factorized manner can substantially improve generative modeling performance from a limited amount of data:
@@
Compositional Visual Generation.  We further consider modeling a probability distribution $p(x\mid T)$ in text-to-image synthesis, where $x$ is an image and $T$ is a complex text description. While this distribution is usually characterized by a single generative model, we can factor the generation as a product of distributions (Liu et al., 2022) given sentences $t_{1}$, $t_{2}$, and $t_{3}$ in the description $T$
@@
Composing Language Models.  Finally, we consider modeling a probability distribution $p(x)$ over a language sequence $x$. Similar to the previous examples, we can represent the likelihood as a composition $p(x)\propto\prod_{i}p_{i}(x)$, where each distribution $p_{i}(x)$ is parameterized by a separate language model. However, directly sampling from such a composition of language models is difficult as it requires intermediate access to the output logits of each model, which are often unavailable for proprietary models. One approach to avoid this issue is to combine outputs of individual language models $p_{i}(x)$ in the language space and use the result as context for representing the final distribution $p(x)$ over language sequences (Du et al., 2023b).
In  Du et al. (2023b), this compositional approach is found to effectively improve the performance of base language models. For instance, on the MATH dataset (Hendrycks et al., 2021), by composing 5 instances of a GPT-3.5 model, we can obtain a final accuracy of $58.0\pm 2.8\%$, even outperforming a much larger and expensive GPT-4 model, which obtains a performance of $55.0\pm 2.9\%$.
@@ 3 Generalization to New Distributions
U-Maze Refer to caption
Medium Refer to caption
Large Refer to caption
@@
corresponding to solving the constraint satisfaction problem. Such an approach enables effective generalization to new problems (Yang et al., 2023b), to temporally extended plans (Mishra et al., 2023), and the combination of heterogenous policies (Wang et al., 2024).
@@
Figure 10: Image Tapestries through Composition. By composing a set of probability distributions defined over different spatial regions in an image, we can construct detailed image tapestries. Figure adapted from (Du et al., 2023a).
@@
where each distribution is defined over bounding boxes in an image. In Figure 10, we illustrate the efficacy of this approach for constructing complex images. This approach enables the synthesis of image tapestries (Du et al., 2023a) and collages (Zhang et al., 2023a).
@@ 4 Generative Modeling with Learned Compositional Structure
where $\text{Enc}(\cdot)$ is a learned neural encoder with low-dimensional latent output to encourage each component to capture distinct regions of an image. By training models to autoencode images with this likelihood expression, each component distribution $p_{\theta}(x\mid\text{Enc}_{i}(x))$ finds interpretable decomposition of images corresponding to individual objects in a scene as well global factors of variation in the scene such as lighting (Du et al., 2021; Su et al., 2024). In Figure 12, we illustrate how these discovered components, $p_{\theta}(x\mid z_{1})$ and $p_{\theta}(x\mid z_{2})$ from a model trained on cubes and spheres, $p_{\phi}(x\mid z_{3})$ and $p_{\phi}(x\mid z_{4})$ from a separate model trained on trucks and boots can be composed together to form the distribution
@@
Discovering Object Classes From Image Distributions.  Given a distribution of images $p(x)$ representing images drawn from different classes in Imagenet, we can model the likelihood of the distribution as a composition
$p_{\theta}(x)\propto p_{\phi}(w\mid x)\prod_{i}p_{\theta}^{i}(x)^{w_{i}},$
where $w_{i}$ refers to the weighting coefficient for each component. In Figure 14, we illustrate that the discovered components in this setting represent each of the original Imagenet classes in the input distribution of images. We further illustrate how these discovered components to be composed together to generate images with multiple classes of objects.
@@ 5.1 Challenges With Sampling from Compositional Distributions
where we need to marginalize over all possible future values of $x_{i+1:T}$. Since this marginalization is different dependent on the value of $x_{i}$, $p_{\text{product}}(x_{i}|x_{0:i-1})$ is not equivalent to $p_{1}(x_{i}|x_{0:i-1})p_{2}(x_{i}|x_{0:i-1})$ and therefore autoregressive factorizations are not directly compositional. Similarly, two learned score functions from diffusion models are not directly composable as they do not correspond to the noisy gradient of the product distribution (Du et al., 2023a).
@@ 5.2 Effective Compositional Sampling on Continuous Distributions
There has been a substantial body of recent work on improving learning in EBMs (Du & Mordatch, 2019; Nijkamp et al., 2019; Grathwohl et al., 2019; Du et al., 2020b; Grathwohl et al., 2021) but EBMs still lag behind other generative approaches in efficiency and scalability of training. By leveraging the close connection of diffusion models with EBMs in  (Song & Ermon, 2019) we can also directly implement the compositional operations with EBMs with diffusion models  (Du et al., 2023a), which we briefly describe below.
Given a diffusion model representing a distribution $p(x)$, we can interpret the $T$ learned denoising functions $\epsilon(x,t)$ of the diffusion model as representing $T$ separate EBM distributions, $e^{-E(x,t)}$, where $\nabla_{x}{E(x,t)}=\epsilon(x,t)$. This sequence of EBM distributions transition from $e^{-E(x,T)}$ representing the Gaussian distribution $\mathcal{N}(0,1)$ to $e^{-E(x,0)}$ representing the target distribution $p_{i}(x)$. We can draw samples from this sequence of EBMs using annealed importance sampling (Du et al., 2023a), where we initialize a sample from Gaussian noise and sequentially run several steps of MCMC on each EBM distribution, starting at $e^{-E(x,T)}$ and ending at $e^{-E(x,0)}$.
This EBM interpretation of diffusion models allows them to be composed using operations such as Equation 2 by applying the operation to each intermediate EBM corresponding to the component diffusion distributions, for instance $e^{-(E_{1}(x,k)+E_{2}(x,k))}$. We can then use an annealed importance sampling procedure on this sequence of composite EBMs. Note that this annealed importance procedure is necessary for accurate compositional sampling – using the reverse diffusion process directly on this composed score does not sample from the composed distribution (Du et al., 2023a).
@@ 5.3 Effective Compositional Sampling on Discrete Distributions
This procedure allows us to leverage $e^{-E_{1}(x)}$ to guide sampling from $e^{-E(x)}$
@@ 6 Discussion and Future Directions
Such compositional systems have additional benefits in terms of both buildability and interpretability. As individual models are responsible for independent subsets of data, each model can be built separately and modularly by different institutions. Simultaneously, at execution time, it is significantly easier to understand and monitor the execution of each simpler constituent model than a single large monolithic model.
In addition, such compositional systems can be more environmentally friendly and easier to deploy than large monolithic models. Since individual models are substantially smaller, they can run efficiently using small amounts of computation. In addition, it is more straightforward to deploy separate models across separate computational machines.
@@
Finally, while we have provided a few promising results on applications of compositional generative modeling, there are many limitations to address in future work. First, the current work on compositional modeling assumes a fixed prespecified structure through which models are composed, limiting generalization to new distributions. To flexibly apply compositional models across new tasks, it would be important to construct systems that can instead automatically discover the correct compositional structure between models as well as the appropriate per-model weighting.
Second, current work on discovering compositional structure assumes that data is naturally factorized into an independent product of components. In many real-world settings, gathered data will often exhibit spurious correlations that violate such independence assumptions, causing existing algorithms to fail to discover the correct structure. Exploring more robust approaches to discovering compositional structure such as through prior knowledge or active intervention in the environment are rich directions for future work.
Lastly, while our focus in this position paper has been on combining separately trained generative models, it would be interesting to theoretically characterize compositional generalization in such systems as well as alternative approaches to improve such generalization. Past theoretical work has characterized compositional generalization in additive models  (Wiedemer et al., 2024; Lachapelle et al., 2024), and it would be interesting to extend such analysis to compositional generative modeling. Furthermore, it would be interesting to explore adding explicit compositional structure to individual models to improve compositional generalization (Misino et al., 2022; Sehgal et al., 2023).
Acknowledgements
We acknowledge support from NSF grant 2214177; from AFOSR grant FA9550-22-1-0249; from ONR MURI grant N00014-22-1-2740; and from ARO grant W911NF-23-1-0034. Yilun Du is supported by a NSF Graduate Fellowship.
Impact Statement
In this paper, we argue that generative models should be built compositionally, from simpler individual parts and illustrate how this enables from data-efficient generative modeling. As generative models become increasingly deployed into production, we believe that such an approach can significantly broaden the impact of such models, enabling them to be deployed in a variety of domains with limited data.